a new car sells for 18000 depreciates 25% a year. what is the value after 4 years?
(this is just an example so i know how to do the rest)

Answer by rfer(16322)

18000*0.75=13500
13500*0.75=10125
10125*0.75=7593.75
7593.75*0.75=$5695.31